OpenStudy (anonymous):

First you will need to create your own rational expression. This will be your “game piece.” You will also need a coin with two different sides to keep track of “heads” versus “tails.”

OpenStudy (anonymous):

Create a rational expression to be your game piece. You may choose from the list of factors below or make your own. There must be a variable term in both the numerator and denominator. (5x) (2x) (x + 4) (x – 5) (2x + 1) (3x + 5)

OpenStudy (anonymous):

lol i had to finish this too

OpenStudy (anonymous):

so I need to create an expression with a variable in the numerator and denominator and the expression needs to be rational...

OpenStudy (anonymous):

@ganeshie8 @phi

OpenStudy (phi):

pick a number between 1 and 100

OpenStudy (anonymous):

4

OpenStudy (phi):

that means you picked the 4th item on your list as the top (numerator) of your rational function so far you have \[ \frac{x-5}{?} \] now pick a different number (or just pick a different item from your list) to get the bottom term

OpenStudy (anonymous):

5x

OpenStudy (phi):

ok, your starting expression is \[ \frac{x-5}{5x} \] now they want you to pick heads or tails for the next step.

OpenStudy (anonymous):

Here is the full list of steps: Create a rational expression to be your game piece. You may choose from the list of factors below or make your own. There must be a variable term in both the numerator and denominator. (5x) (2x) (x + 4) (x – 5) (2x + 1) (3x + 5) Turn one. Flip your coin and perform the appropriate operation. Explain to the game master how to add your rational expression to the one on the correct space. Use complete sentences. Turn two. Flip your coin and perform the appropriate operation. Discuss and identify any possible restrictions that exist with (or in) the resulting rational expression. Turn three. Flip your coin and perform the appropriate operation. Explain to the game master how to multiply your rational expression to the one on the correct space. Use complete sentences. Turn four. Flip your coin and perform the appropriate operation. Discuss why the degree of the resulting denominator did not change from your expression’s degree. Turn five, the final level! Perform the appropriate operation. Using complete sentences, describe the steps you used.

OpenStudy (anonymous):

Thats all the questions I have to answer for this

OpenStudy (phi):

yes, I got the idea. And the next step for you is flip a coin.

OpenStudy (anonymous):

Heads :D

OpenStudy (phi):

congratulations. You just took the first steps to finishing this problem. they want you to add \[ \frac{x-5}{5x}+\frac{x+1}{2x} \]

OpenStudy (phi):

the question is a bit dubious, because that is the answer, as I posted it. However, I am sure they want you to put it over a common denominator and simplify as much as possible.

OpenStudy (phi):

both denominators have an x , so you are good there. But you need a common denominator for the 2 and 5

OpenStudy (anonymous):

How about x - 10/ / 10x + x + 5 / 10x?

OpenStudy (phi):

close. but the idea is to multiply the top and bottom of the first fraction by 2 \[ \frac{2}{2} \] that gives you 10x in the bottom but when you multiply the top by 2, that means you multiply *everything* up top by 2 ditto for the second fraction (except you use 5/5 )

OpenStudy (anonymous):

\[\frac{ x^2 - 10 }{ 10x } \frac{ x^2 + 5 }{ 10x }\]